If you are 30 m from the foot of the Leaning Tower of Pisa, the top of the Leaning Tower is at 65 degree angle of of elevation. If the tower were perpendicular to the ground, the tower is inclined towards us 4” . Calculate to find the height of the tower with such data?


1
Expert's answer
2022-03-24T18:56:14-0400

Solution

To find the height of the tower we use sine rule which states

"\frac{a}{sineA}=\frac{b}{sineB}"


"\frac{30}{sine4°}=\frac{b}{sine111°}"


cross multiplication


"b=\frac{30sine111°}{sine4°}"


b=34.27m